# Deep Learning for 3D Human Pose Estimation and Mesh Recovery: A Survey

### v1

@@ 4.2 Multi-person 3D pose estimation
As depicted in Fig.5, multi-person 3D pose estimation can be divided into two main categories: the bottom-up method (estimation + association) and the top-down method (detection + estimation). With a two-stage pipeline, the top-down method first detects every person in the input images and then extracts each person’s keypoints in the previously detected bounding box. The bottom-up method initially detects all keypoints in one stage and subsequently associates them with their respective persons. For instance, Cheng et al. cheng2022dual integrated both top-down and bottom-up approaches in multi-person pose estimation, complementing each other’s shortcomings. The top-down method shows robustness against potential erroneous bounding boxes, while the bottom-up network is more robust in handling scale variations. Finally, the 3D poses estimated from both top-down and bottom-up networks are fed into an integrated network to obtain the final 3D pose. Some single-stage methods and approaches combine the two methods as mentioned above. Unlike mainstream two-stage solutions for multi-person 3D human pose estimation, Jin et al. jin2022single introduced a single-stage method. This method expresses the 2D pose in the image plane and the depth information of a 3D body instance via the proposed decoupled representation and predicts the scale information of instances by extracting 2D pose features and enabling depth regression.
@@ 4.2.1 Top-down methods
In top-down methods, AlphaPose fang2022alphapose predicts whole-body multi-person 3D human pose including face, body, hand, and foot. In the proposed framework, the symmetric integral keypoint regression module achieves fast and fine localization, the parametric pose non-maximum suppression module helps to eliminate redundant human detections, and the pose aware identity embedding module enables joint pose estimation and tracking. To achieve better real-time performance in multi-person 3D pose estimation, Chen et al. chen2020cross proposed a multi-view temporal consistency method in real-time from videos, which matches the 2D inputs with 3D poses directly in three-dimensional space and achieves over 150 FPS (frames per second) on a 12-camera setup and 34 FPS on a 28-camera setup. Additionally, choosing appropriate representations can effectively improve multi-person pose estimation results. For instance, Zhang et al. zhang2022voxeltrack employed a voxel representation to predict multi-person 3D pose and tracking, where the proposed voxel representation can determine whether each voxel contains a particular body joint. In multi-person scenarios, occlusion issues may be more severe and complex than in single-person scenarios. To address this challenge, Wu et al. wu2021graph utilized graph neural networks to boost information passing efficiency for multi-person, multi-view 3D human pose estimation. The multi-view matching graph module associates coarse cross-view poses within the networks, and the center refinement graph module further refines the results. Single-shot learning trains the model using significantly less data than is typically required for supervised learning, aiming to improve effectiveness. PandaNet benzine2020pandanet as an anchor-based model introduces a pose-aware anchor selection strategy to discard ambiguous anchors. Moon et al. moon2019camera proposed a top-down, camera distance-aware method for multi-person 3D pose estimation without relying on ground-truth information.
@@ 4.2.2 Bottom-up methods
MubyNet zanfir2018deep is a typical bottom-up, multi-task method for multi-person 3D pose estimation, which allows for training all component parameters. In the research on the lightweight top-down framework, Fabbri et al. fabbri2020compressed utilized high-resolution volumetric heatmaps to improve the estimation performance and designed a volumetric heatmap autoencoder that can compress the size of the representation. Designing better supervisory methods could also be effective, Wang et al. wang2020hmor introduced a novel supervisory approach named Hierarchical Multi-person Ordinal Relations (HMOR) using a monocular camera and designed a comprehensive top-level model to learn these ordinal relations, enhancing the accuracy of human depth estimation through a coarse-to-fine architecture. In the top-down estimation on single-shot learning, Mehta et al. mehta2018single inferred whole body pose under strong partial occlusions through occlusion-robust pose-maps in their proposed single-shot framework. Zhen et al. zhen2020smap designed a single-shot, bottom-up framework that estimates the absolute positions of multiple people by leveraging depth-related cues across the entire image. After their previous work benzine2020pandanet , Benzine et al. benzine2021single proposed a single-shot 3D human pose estimation method that predicts multi-person 3D poses without the need for bounding boxes. This method extends the Stacked Hourglass Network newell2016stacked to handle multi-person situations. To address the issue of occlusion, LCR-Net++ rogez2019lcr integrates adjacent pose hypotheses to predict the multi-person 2D and 3D poses without approximating initial human localization when a person is partially occluded or truncated by the image boundaries.
@@ 4.3 Summary of 3D pose estimation
Human mesh recovery can be divided into two categories based on their representation models: template-based (parametric) methods and template-free (non-parametric) methods, as shown in Fig.6. Template-based human mesh recovery reconstructs predefined models (such as SCAPE anguelov2005scape , SMPL loper2015smpl ) by estimating the model’s parameters. In contrast, template-free human mesh recovery predicts the 3D body directly from input data without reliance on predefined models. The parametric approaches can be more robust than non-parametric methods with the templates’ prior knowledge, but their flexibility and detail are inherently limited even when extended with more parameters like SMPL+H romero2022embodied and SMPL+X cai2023smpler . Human mesh recovery is a crucial technique for digitizing the human body, posing significant challenges in computer vision and computer graphics due to the complex nature of geometric textures and color variations. Moreover, human mesh recovery faces challenges similar to 3D pose estimation, including environmental interference, multi-person scenarios, and occlusion issues. In this section, we will introduce the dominant methods based on these categories and challenges, with a comprehensive summary provided in Table 2.
@@
Template-based Naked Multimodal Methods $\bullet$ Hybrid annotations: Rong et al. rong2019delving
$\bullet$ Optical flow: DTS-VIBE li2022deep
$\bullet$ Silhouettes: LASOR yang2022lasor
$\bullet$ Cropped image and bounding box: CLIFF li2022cliff
Utilizing Attention Mechanism $\bullet$ Part-driven attention: PARE kocabas2021pare
$\bullet$ Graph attention: Mesh Graphormer lin2021mesh
$\bullet$ Spatio-temporal attention: MPS-Net wei2022capturing , PSVT qiu2023psvt
$\bullet$ Efficient architecture: FastMETRO cho2022cross , Xue et al. xue20223d
$\bullet$ End-to-end structure: METRO lin2021end
@@
Exploiting Temporal Information $\bullet$ Temporally encoding features: Kanazawa et al. kanazawa2019learning
$\bullet$ Self-attention temporal: VIBE kocabas2020vibe
$\bullet$ Temporally consistent: TCMR choi2021beyond
$\bullet$ Multi-level spatial-temporal attention: MAED wan2021encoder
$\bullet$ Temporally embedded live stream: TePose wang2022live
$\bullet$ Short-term and long-term temporal correlations: GLoT shen2023global
@@
Multi-view Methods $\bullet$ Confidence-aware majority voting mechanism: Dong et al. dong2021shape
$\bullet$ Probabilistic-based multi-view: Sengupta et al. sengupta2021probabilistic
$\bullet$ Dynamic physics-geometry consistency: Huang et al. huang2021dynamic
$\bullet$ Cross-view fusion: Zhuo et al. zhuo2023towards
Boosting Efficiency $\bullet$ Sparse constrained formulation: SCOPE fan2021revitalizing
$\bullet$ Single-stage model: BMP zhang2021body
$\bullet$ Process heatmap inputs: HeatER zheng2022heater
$\bullet$ Removing redundant tokens: TORE dou2023tore
Developing Various Representations $\bullet$ Texture map: TexturePose pavlakos2019texturepose
$\bullet$ UV map: Zhang et al. zhang2020learning , DecoMR zeng20203d , Zhang et al. zhang2020object
$\bullet$ Heat map: Sun et al. sun2021monocular , 3DCrowdNet choi2022learning
$\bullet$ Uniform representation: DSTformer zhu2023motionbert
Utilizing Structural Information $\bullet$ Part-based: holopose guler2019holopose
$\bullet$ Skeleton disentangling: Sun et al. sun2019human
$\bullet$ Hybrid inverse kinematics: HybrIK li2021hybrik , NIKI li2023niki
$\bullet$ Uncertainty-aware: Lee et al. lee2021uncertainty
$\bullet$ Kinematic tree structure: Sengupta et al. sengupta2021hierarchical
$\bullet$ Kinematic chains: SGRE wang20233d
@@
Choosing Appropriate Learning Strategies $\bullet$ Self-improving: SPIN kolotouros2019learning , ReFit wang2023refit , You et al. you2023co
$\bullet$ Novel losses: Zanfir et al. zanfir2020weakly , Jiang et al. jiang2020coherent
$\bullet$ Unsupervised learning: Madadi et al. madadi2021deep , Yu et al. yu2021skeleton2mesh
$\bullet$ Bilevel online adaptation: Guan et al. guan2022out
$\bullet$ Single-shot: Pose2UV huang2022pose2uv
$\bullet$ Contrastive learning: JOTR li2023jotr
$\bullet$ Domain adaptation: Nam et al. nam2023cyclic
Detailed With Clothes $\bullet$ Alldieck et al. alldieck2019learning
$\bullet$ Multi-Garment Network (MGN) bhatnagar2019multi
$\bullet$ Texture map: Tex2Shape alldieck2019tex2shape
$\bullet$ Layered garment representation: BCNet jiang2020bcnet
$\bullet$ Temporal span: H4D jiang2022h4d
With Hands $\bullet$ Linguistic priors: SGNify forte2023reconstructing
$\bullet$ Two-hands interaction: zhang2021interacting
$\bullet$ Hand-object interaction: chen2021joint
Whole Body $\bullet$ PROX hassan2019resolving
$\bullet$ ExPose choutas2020monocular
$\bullet$ FrankMocap rong2021frankmocap
$\bullet$ PIXIE feng2021collaborative
$\bullet$ Moon et al. moon2022accurate
$\bullet$ PyMAF zhang2023pymaf
$\bullet$ OSX lin2023one
$\bullet$ HybrIK-X li2023hybrik
Template-free Regression-based $\bullet$ FACSIMILE smith2019facsimile , PeeledHuman jinka2020peeledhuman , GTA zhang2023global , NSF xue2023nsf
Optimization-based Differentiable $\bullet$ DiffPhy gartner2022differentiable , AG3D dong2023ag3d
Implicit Representations $\bullet$ PIFu saito2019pifu , PIFuHD saito2020pifuhd
$\bullet$ Canonical space: ARCH huang2020arch , ARCH++ he2021arch++ , CAR liao2023high
$\bullet$ Geometric priors: GeoPIFu he2020geo
$\bullet$ Novel representations: Peng et al. peng2021neural , 3DNBF zhang20233d
Neural Radiance Fields $\bullet$ Volume deformation scheme gao2022mps
$\bullet$ ActorsNeRF mu2023actorsnerf
Diffusion Models $\bullet$ HMDiff foo2023distribution
@@
Implicit + Explicit $\bullet$ HMD zhu2019detailed , IP-Net bhatnagar2020combining , PaMIR zheng2021pamir , Zhu et al. zhu2021detailed , ICON xiu2022icon , ECON xiu2023econ , DELTA feng2023learning , GETAvatar zhang2023getavatar
Diffusion + Explicit $\bullet$ DINAR svitov2023dinar
NeRF + Explicit $\bullet$ TransHuman pan2023transhuman
Gaussian Splatting + Explicit $\bullet$ Animatable 3D Gaussian liu2023animatable
@@ 5.1.1 Naked human body recovery
Multimodal Methods. Multimodality in human mesh recovery harnesses the potential by combining various modalities of data, such as RGB images, depth information, and optical flow. Integrating data in different modalities can significantly enhance the robustness and precision of mesh recovery. Rong et al. rong2019delving proposed a hybrid annotation method with a hybrid training strategy for human mesh recovery to reduce annotation costs, which effectively utilizes various types of heterogeneous annotations, including 3D and 2D annotations, body part segmentation, and dense correspondence. Deep Two-Stream Video Inference for Human Body Pose and Shape Estimation (DTS-VIBE) li2022deep method redefines the task as a multimodal problem, merging RGB data with optical flow to achieve a more reliable estimation and address temporal inconsistencies from RGB videos. LASOR yang2022lasor estimates 3D pose and shape by synthesizing occlusion-aware silhouettes and 2D keypoints data in scenes with inter-person occlusion. CLIFF li2022cliff utilizes cropped images and bounding box information from the pre-training phase as inputs to enhance the accuracy of global rotation estimation in the camera coordinate system.
@@
Utilizing Attention Mechanism. Transformer vaswani2017attention as a self-attention model has demonstrated remarkable success in Natural Language Processing (NLP) devlin2018bert brown2020language . Following this, the Vision Transformer (ViT) dosovitskiy2020image has mirrored these successes in the field of computer vision tolstikhin2021mlp liu2021swin , and numerous Transformer-based methods are now being employed in mesh recovery. The attention mechanism serves to amplify the importance of certain parts in the neural network. To address the partial occlusion issues, Part Attention REgressor (PARE) kocabas2021pare leverages the relationships between body parts derived from segmentation masks, prompting the network to enhance predictions for occluded body parts. Mesh Graphormer lin2021mesh utilizes a GCNN-reinforced transformer for estimating 3D mesh vertices and body joints, while a GCNN is employed to infer interactions among neighboring vertices based on pre-existing mesh topology. The architecture effectively merges graph-based networking with the attention mechanism of transformers, enabling the modeling of local and global interactions. MPS-Net wei2022capturing employs motion continuity attention to capture temporal coherence. This approach then leverages a hierarchical attentive feature mechanism to combine features from temporally adjacent representations more effectively. FastMETRO cho2022cross leverages self-attention mechanisms for non-adjacent vertices based on the topology of the body’s triangular mesh to minimize memory overhead and accelerate inference speed. Xue et al. xue20223d proposed a learnable sampling module for human mesh recovery to reduce inherent depth ambiguity. This module aggregates global information by generating joint adaptive tokens, utilizing non-local information within the input image. METRO lin2021end is a mesh transformer for end-to-end human mesh recovery, in which the encoder captures interactions between vertices and joints, and the decoder outputs 3D joint coordinates and the mesh structure. Qiu et al. qiu2023psvt proposed an end-to-end Transformer-based method for multi-person human mesh recovery in videos. This approach utilizes a spatio-temporal encoder to extract global features, followed by a spatio-temporal pose and shape decoder to predict human pose and mesh.
@@
Exploiting Temporal Information. With the advancement of video technology, video-based human mesh recovery methods that extract temporal information from adjacent frames have shown increased potential. To learn the 3D dynamics of the human body more accurately and effectively, Kanazawa et al. kanazawa2019learning proposed a framework that produces smooth 3D meshes from videos. This framework also predicts past and future 3D motions by temporally encoding image features. Video Inference for Body pose and shape Estimation (VIBE) kocabas2020vibe estimates kinematically plausible motion sequences through self-attention mechanism-based temporal network and adversarial training without requiring any ground-truth 3D labels. To diminish the dependency on static features in the current frame, addressing a limitation of previous temporal-based methods, Choi et al. choi2021beyond developed the Temporally Consistent Mesh Recovery (TCMR) system. The system utilizes temporal information to ensure consistency and effectively recovers smooth 3D human motion by incorporating data from past and future frames. Multi-level Attention Encoder-Decoder (MAED) network wan2021encoder captures relationships at multiple levels, including the spatial-temporal and human joint levels, through its multi-level attention mechanism. Wang et al. wang2022live introduced the Temporally embedded 3D human body Pose and shape estimation (TePose) method, specifically tailored for live stream videos. They designed a motion discriminator for adversarial training, utilizing datasets without any 3D labels, through a multi-scale spatio-temporal graph convolutional network. Additionally, they employed a sequential data loading strategy to accommodate the unique start-to-end data processing requirements of live streaming. To effectively balance the learning of short-term and long-term temporal correlations, Global-to-Local Transformer (GLoT) shen2023global structurally decouples the modeling of long-term and short-term correlations.
@@
Multi-view Methods. Dong et al. dong2021shape designed a practical multi-view framework that combines 2D observations from multi-view images into a unified 3D representation for individual instances using a confidence-aware majority voting mechanism. Sengupta et al. sengupta2021probabilistic introduced a probabilistic-based multi-view method without constraints such as specific target poses, viewpoints, or background conditions across image sequences. Huang et al. huang2021dynamic proposed a dynamic physics-geometry consistency approach for multi-person multi-view mesh recovery. This method integrates motion priors, extrinsic camera parameters, and human mesh data to mitigate the impact of noisy human semantic data. Zhuo et al. zhuo2023towards proposed a cross-view fusion method that predicts foot posture by achieving a more refined 3D intermediate representation and alleviating inconsistencies across different views.
@@
Boosting Efficiency. Maintaining excellent performance with a lightweight model and low computational cost is essential and challenging, especially in applications such as wearable devices, power-limited systems, and edge computing. SCOPE fan2021revitalizing efficiently computes the Gauss-Newton direction using 2D and 3D keypoints for human mesh recovery. The method capitalizes on inherent sparsity and employs a sparse constrained formulation, achieving real-time performance at over 30 FPS. To implement a single-stage multi-person human mesh recovery model, Body Meshes as Points (BMP) zhang2021body represents multiple people as points and associates each with a single body mesh to significantly improve efficiency and performance. HeatER zheng2022heater processes heatmap inputs directly to reduce memory and computational costs. Dou et al. dou2023tore designed an efficient transformer for human mesh recovery to reduce model complexity and computational cost by removing redundant tokens.

### v2

@@ 4.2 Multi-person 3D pose estimation
As depicted in Fig.6, multi-person 3D pose estimation can be divided into two main categories: the bottom-up method (estimation + association) and the top-down method (detection + estimation). With a two-stage pipeline, the top-down method first detects every person in the input images and then extracts each person’s keypoints in the previously detected bounding box. The bottom-up method initially detects all keypoints in one stage and subsequently associates them with their respective persons. For instance, Cheng et al. [121] integrated both top-down and bottom-up approaches in multi-person pose estimation, complementing each other’s shortcomings. The top-down method shows robustness against potential erroneous bounding boxes, while the bottom-up network is more robust in handling scale variations. Finally, the 3D poses estimated from both top-down and bottom-up networks are fed into an integrated network to obtain the final 3D pose. Some single-stage methods and approaches combine the two methods as mentioned above. Unlike mainstream two-stage solutions for multi-person 3D human pose estimation, Jin et al. [120] introduced a single-stage method. This method expresses the 2D pose in the image plane and the depth information of a 3D body instance via the proposed decoupled representation and predicts the scale information of instances by extracting 2D pose features and enabling depth regression.
@@ 4.2.1 Top-down methods
In top-down methods, AlphaPose [111] predicts whole-body multi-person 3D human pose including face, body, hand, and foot. In the proposed framework, the symmetric integral keypoint regression module achieves fast and fine localization, the parametric pose non-maximum suppression module helps to eliminate redundant human detections, and the pose aware identity embedding module enables joint pose estimation and tracking. To achieve better real-time performance in multi-person 3D pose estimation, Chen et al. [110] proposed a multi-view temporal consistency method in real-time from videos, which matches the 2D inputs with 3D poses directly in three-dimensional space and achieves over 150 FPS (frames per second) on a 12-camera setup and 34 FPS on a 28-camera setup. Additionally, choosing appropriate representations can effectively improve multi-person pose estimation results. For instance, Zhang et al. [2] employed a voxel representation to predict multi-person 3D pose and tracking, where the proposed voxel representation can determine whether each voxel contains a particular body joint. In multi-person scenarios, occlusion issues may be more severe and complex than in single-person scenarios. To address this challenge, Wu et al. [112] utilized graph neural networks to boost information passing efficiency for multi-person, multi-view 3D human pose estimation. The multi-view matching graph module associates coarse cross-view poses within the networks, and the center refinement graph module further refines the results. Single-shot learning trains the model using significantly less data than is typically required for supervised learning, aiming to improve effectiveness. PandaNet [52] as an anchor-based model introduces a pose-aware anchor selection strategy to discard ambiguous anchors. Moon et al. [113] proposed a top-down, camera distance-aware method for multi-person 3D pose estimation without relying on ground-truth information.
@@ 4.2.2 Bottom-up methods
MubyNet [124] is a typical bottom-up, multi-task method for multi-person 3D pose estimation, which allows for training all component parameters. In the research on the lightweight top-down framework, Fabbri et al. [114] utilized high-resolution volumetric heatmaps to improve the estimation performance and designed a volumetric heatmap autoencoder that can compress the size of the representation. Designing better supervisory methods could also be effective, Wang et al. [115] introduced a novel supervisory approach named Hierarchical Multi-person Ordinal Relations (HMOR) using a monocular camera and designed a comprehensive top-level model to learn these ordinal relations, enhancing the accuracy of human depth estimation through a coarse-to-fine architecture. In the top-down estimation on single-shot learning, Mehta et al. [118] inferred whole body pose under strong partial occlusions through occlusion-robust pose-maps in their proposed single-shot framework. Zhen et al. [116] designed a single-shot, bottom-up framework that estimates the absolute positions of multiple people by leveraging depth-related cues across the entire image. After their previous work [52], Benzine et al. [117] proposed a single-shot 3D human pose estimation method that predicts multi-person 3D poses without the need for bounding boxes. This method extends the Stacked Hourglass Network [125] to handle multi-person situations. To address the issue of occlusion, LCR-Net++ [119] integrates adjacent pose hypotheses to predict the multi-person 2D and 3D poses without approximating initial human localization when a person is partially occluded or truncated by the image boundaries.
@@ 4.3 Summary of 3D pose estimation
Human mesh recovery can be divided into two categories based on their representation models: template-based (parametric) methods and template-free (non-parametric) methods, as shown in Fig.7. Template-based human mesh recovery reconstructs predefined models (such as SCAPE [34], SMPL [35]) by estimating the model’s parameters. In contrast, template-free human mesh recovery predicts the 3D body directly from input data without reliance on predefined models. The parametric approaches can be more robust than non-parametric methods with the templates’ prior knowledge, but their flexibility and detail are inherently limited even when extended with more parameters like SMPL+H [36] and SMPL+X [61]. Human mesh recovery is a crucial technique for digitizing the human body, posing significant challenges in computer vision and computer graphics due to the complex nature of geometric textures and color variations. Moreover, human mesh recovery faces challenges similar to 3D pose estimation, including environmental interference, multi-person scenarios, and occlusion issues. In this section, we will introduce the dominant methods based on these categories and challenges, with a comprehensive summary provided in Table 2.
@@
Template-based Naked Multimodal Methods $\bullet$ Hybrid annotations: Rong et al. [126]
$\bullet$ Optical flow: DTS-VIBE [127]
$\bullet$ Silhouettes: LASOR [128]
$\bullet$ Cropped image and bounding box: CLIFF [129]
Utilizing Attention Mechanism $\bullet$ Part-driven attention: PARE [130]
$\bullet$ Graph attention: Mesh Graphormer [131]
$\bullet$ Spatio-temporal attention: MPS-Net [132], PSVT [133]
$\bullet$ Efficient architecture: FastMETRO [134], Xue et al. [135]
$\bullet$ End-to-end structure: METRO [136]
@@
Exploiting Temporal Information $\bullet$ Temporally encoding features: Kanazawa et al. [137]
$\bullet$ Self-attention temporal: VIBE [138]
$\bullet$ Temporally consistent: TCMR [139]
$\bullet$ Multi-level spatial-temporal attention: MAED [140]
$\bullet$ Temporally embedded live stream: TePose [141]
$\bullet$ Short-term and long-term temporal correlations: GLoT [142]
@@
Multi-view Methods $\bullet$ Confidence-aware majority voting mechanism: Dong et al. [143]
$\bullet$ Probabilistic-based multi-view: Sengupta et al. [144]
$\bullet$ Dynamic physics-geometry consistency: Huang et al. [33]
$\bullet$ Cross-view fusion: Zhuo et al. [145]
Boosting Efficiency $\bullet$ Sparse constrained formulation: SCOPE [146]
$\bullet$ Single-stage model: BMP [147]
$\bullet$ Process heatmap inputs: HeatER [148]
$\bullet$ Removing redundant tokens: TORE [149]
Developing Various Representations $\bullet$ Texture map: TexturePose [150]
$\bullet$ UV map: Zhang et al. [151], DecoMR [152], Zhang et al. [153]
$\bullet$ Heat map: Sun et al. [154], 3DCrowdNet [155]
$\bullet$ Uniform representation: DSTformer [156]
Utilizing Structural Information $\bullet$ Part-based: holopose [157]
$\bullet$ Skeleton disentangling: Sun et al. [158]
$\bullet$ Hybrid inverse kinematics: HybrIK [159], NIKI [160]
$\bullet$ Uncertainty-aware: Lee et al. [161]
$\bullet$ Kinematic tree structure: Sengupta et al. [162]
$\bullet$ Kinematic chains: SGRE [163]
@@
Choosing Appropriate Learning Strategies $\bullet$ Self-improving: SPIN [164], ReFit [165], You et al. [5]
$\bullet$ Novel losses: Zanfir et al. [48], Jiang et al. [166]
$\bullet$ Unsupervised learning: Madadi et al. [167], Yu et al. [50]
$\bullet$ Bilevel online adaptation: Guan et al. [168]
$\bullet$ Single-shot: Pose2UV [169]
$\bullet$ Contrastive learning: JOTR [170]
$\bullet$ Domain adaptation: Nam et al. [171]
Detailed With Clothes $\bullet$ Alldieck et al. [172]
$\bullet$ Multi-Garment Network (MGN) [173]
$\bullet$ Texture map: Tex2Shape [174]
$\bullet$ Layered garment representation: BCNet [175]
$\bullet$ Temporal span: H4D [39]
With Hands $\bullet$ Linguistic priors: SGNify [176]
$\bullet$ Two-hands interaction: [177]
$\bullet$ Hand-object interaction: [178]
Whole Body $\bullet$ PROX [179]
$\bullet$ ExPose [180]
$\bullet$ FrankMocap [181]
$\bullet$ PIXIE [182]
$\bullet$ Moon et al. [183]
$\bullet$ PyMAF [184]
$\bullet$ OSX [185]
$\bullet$ HybrIK-X [186]
Template-free Regression-based $\bullet$ FACSIMILE [187], PeeledHuman [188], GTA [189], NSF [190]
Optimization-based Differentiable $\bullet$ DiffPhy [191], AG3D [192]
Implicit Representations $\bullet$ PIFu [193], PIFuHD [194]
$\bullet$ Canonical space: ARCH [195], ARCH++ [196], CAR [197]
$\bullet$ Geometric priors: GeoPIFu [198]
$\bullet$ Novel representations: Peng et al. [199], 3DNBF [200]
Neural Radiance Fields $\bullet$ Volume deformation scheme [201]
$\bullet$ ActorsNeRF [51]
Diffusion Models $\bullet$ HMDiff [202]
@@
Implicit + Explicit $\bullet$ HMD [203], IP-Net [204], PaMIR [42], Zhu et al. [205], ICON [206], ECON [207], DELTA [10], GETAvatar [208]
Diffusion + Explicit $\bullet$ DINAR [209]
NeRF + Explicit $\bullet$ TransHuman [210]
Gaussian Splatting + Explicit $\bullet$ Animatable 3D Gaussian [211]
@@ 5.1.1 Naked human body recovery
Multimodal Methods. Multimodality in human mesh recovery harnesses the potential by combining various modalities of data, such as RGB images, depth information, and optical flow. Integrating data in different modalities can significantly enhance the robustness and precision of mesh recovery. Rong et al. [126] proposed a hybrid annotation method with a hybrid training strategy for human mesh recovery to reduce annotation costs, which effectively utilizes various types of heterogeneous annotations, including 3D and 2D annotations, body part segmentation, and dense correspondence. Deep Two-Stream Video Inference for Human Body Pose and Shape Estimation (DTS-VIBE) [127] method redefines the task as a multimodal problem, merging RGB data with optical flow to achieve a more reliable estimation and address temporal inconsistencies from RGB videos. LASOR [128] estimates 3D pose and shape by synthesizing occlusion-aware silhouettes and 2D keypoints data in scenes with inter-person occlusion. CLIFF [129] utilizes cropped images and bounding box information from the pre-training phase as inputs to enhance the accuracy of global rotation estimation in the camera coordinate system.
@@
Utilizing Attention Mechanism. Transformer [212] as a self-attention model has demonstrated remarkable success in Natural Language Processing (NLP) [213] [214]. Following this, the Vision Transformer (ViT) [46] has mirrored these successes in the field of computer vision [45] [215], and numerous Transformer-based methods are now being employed in mesh recovery. The attention mechanism serves to amplify the importance of certain parts in the neural network. To address the partial occlusion issues, Part Attention REgressor (PARE) [130] leverages the relationships between body parts derived from segmentation masks, prompting the network to enhance predictions for occluded body parts. Mesh Graphormer [131] utilizes a GCNN-reinforced transformer for estimating 3D mesh vertices and body joints, while a GCNN is employed to infer interactions among neighboring vertices based on pre-existing mesh topology. The architecture effectively merges graph-based networking with the attention mechanism of transformers, enabling the modeling of local and global interactions. MPS-Net [132] employs motion continuity attention to capture temporal coherence. This approach then leverages a hierarchical attentive feature mechanism to combine features from temporally adjacent representations more effectively. FastMETRO [134] leverages self-attention mechanisms for non-adjacent vertices based on the topology of the body’s triangular mesh to minimize memory overhead and accelerate inference speed. Xue et al. [135] proposed a learnable sampling module for human mesh recovery to reduce inherent depth ambiguity. This module aggregates global information by generating joint adaptive tokens, utilizing non-local information within the input image. METRO [136] is a mesh transformer for end-to-end human mesh recovery, in which the encoder captures interactions between vertices and joints, and the decoder outputs 3D joint coordinates and the mesh structure. Qiu et al. [133] proposed an end-to-end Transformer-based method for multi-person human mesh recovery in videos. This approach utilizes a spatio-temporal encoder to extract global features, followed by a spatio-temporal pose and shape decoder to predict human pose and mesh.
@@
Exploiting Temporal Information. With the advancement of video technology, video-based human mesh recovery methods that extract temporal information from adjacent frames have shown increased potential. To learn the 3D dynamics of the human body more accurately and effectively, Kanazawa et al. [137] proposed a framework that produces smooth 3D meshes from videos. This framework also predicts past and future 3D motions by temporally encoding image features. Video Inference for Body pose and shape Estimation (VIBE) [138] estimates kinematically plausible motion sequences through self-attention mechanism-based temporal network and adversarial training without requiring any ground-truth 3D labels. To diminish the dependency on static features in the current frame, addressing a limitation of previous temporal-based methods, Choi et al. [139] developed the Temporally Consistent Mesh Recovery (TCMR) system. The system utilizes temporal information to ensure consistency and effectively recovers smooth 3D human motion by incorporating data from past and future frames. Multi-level Attention Encoder-Decoder (MAED) network [140] captures relationships at multiple levels, including the spatial-temporal and human joint levels, through its multi-level attention mechanism. Wang et al. [141] introduced the Temporally embedded 3D human body Pose and shape estimation (TePose) method, specifically tailored for live stream videos. They designed a motion discriminator for adversarial training, utilizing datasets without any 3D labels, through a multi-scale spatio-temporal graph convolutional network. Additionally, they employed a sequential data loading strategy to accommodate the unique start-to-end data processing requirements of live streaming. To effectively balance the learning of short-term and long-term temporal correlations, Global-to-Local Transformer (GLoT) [142] structurally decouples the modeling of long-term and short-term correlations.
@@
Multi-view Methods. Dong et al. [143] designed a practical multi-view framework that combines 2D observations from multi-view images into a unified 3D representation for individual instances using a confidence-aware majority voting mechanism. Sengupta et al. [144] introduced a probabilistic-based multi-view method without constraints such as specific target poses, viewpoints, or background conditions across image sequences. Huang et al. [33] proposed a dynamic physics-geometry consistency approach for multi-person multi-view mesh recovery. This method integrates motion priors, extrinsic camera parameters, and human mesh data to mitigate the impact of noisy human semantic data. Zhuo et al. [145] proposed a cross-view fusion method that predicts foot posture by achieving a more refined 3D intermediate representation and alleviating inconsistencies across different views.
@@
Boosting Efficiency. Maintaining excellent performance with a lightweight model and low computational cost is essential and challenging, especially in applications such as wearable devices, power-limited systems, and edge computing. SCOPE [146] efficiently computes the Gauss-Newton direction using 2D and 3D keypoints for human mesh recovery. The method capitalizes on inherent sparsity and employs a sparse constrained formulation, achieving real-time performance at over 30 FPS. To implement a single-stage multi-person human mesh recovery model, Body Meshes as Points (BMP) [147] represents multiple people as points and associates each with a single body mesh to significantly improve efficiency and performance. HeatER [148] processes heatmap inputs directly to reduce memory and computational costs. Dou et al. [149] designed an efficient transformer for human mesh recovery to reduce model complexity and computational cost by removing redundant tokens.